The standard free energy change, AGo, for the reaction C2H5OH(g) C2H4(g) + H20(g) is +8.25 kJ/mol at 25°C. Calculate Kp for the equilibrium.
Answer #1

T= 25.0 oC

= (25.0+273) K

= 298 K

ΔGo = 8.25 KJ/mol

ΔGo = 8250 J/mol

use:

ΔGo = -R*T*ln Kc

8250 = - 8.314*298.0* ln(Kc)

ln Kc = -3.3299

Kc = 3.58*10^-2

T = 298 K

Δ n = number of gaseous molecule in product - number of gaseous molecule in reactant

Δ n = 1

use:

Kp= Kc (RT)^Δ n

Kp = 3.58*10^-2*(0.08206*298.0)^(1)

Kp = 0.8754

Answer: 0.875